OpenStudy (anonymous):

The cut out lawn mowing company consists of two people: Tina and Bill. IF Tina cuts a lawn by herself she can do it in 4 hrs. If Bill cuts it it takes him an hour longer. How long would it take if they did it together?

OpenStudy (wolf1728):

There is a formula for this Total Time = (Tina * Bill) / (Tina + Bill) Total Time = (4*5) / (4+5) Total Time = 20/9 hours Still you might want a better explanation than that.

OpenStudy (wolf1728):

Tina's rate = .25/hour (meaning .25 of the lawn is cut per hour). Bill's rate = .20/hour So, in one hour, (.25 + .20) part of the lawn is cut. or .45 parts per hour Rate for both Tina and Bill = .45 lawn/one hour We need to know the time for the total lawn so the total time would be hours/lawn. To get hours per lawn, we take the reciprocal of .45/1 hour OR 1/.45 which equals 2.222222 hours (which equals 20/9, my previous answer).